On May 1, 2020, Rolly Industries begins liquidation activities and adopts the liquidation basis of accounting. The book value of its reported assets total $700,000, including $10,000 in cash, and the book value of its liabilities, consisting of bank loans, total $600,000. Expected proceeds from reported assets other than cash are:

  • Receivables, $50,000
  • Inventories, $150,000
  • Plant and equipment, $300,000

Previously unreported identifiable intangible assets have a fair value of $80,000. Expected costs of liquidating assets are $20,000, and negotiations are in process to reduce Rolly’s bank loans by 25%.

During the two months ending June 30, 2020, the following transactions occur:

  • Receivables of $48,000 are collected and the rest are determined to be uncollectible.
  • Inventories are sold for $100,000.
  • Plant and equipment is sold for $125,000.
  • The identifiable intangible assets are sold for $72,000.
  • Liquidation costs of $10,000 are paid.
  • Bank loans of $325,000 are paid, and creditors holding $275,000 of loans agree to accept $250,000 as full payment.
  • Fair values of remaining assets other than cash are:
  • Inventories, $55,000
  • Plant and equipment, $185,000
  • Estimated future liquidation costs are $6,000

1. On the statement of net assets in liquidation at June 30, 2020, total assets are:

a.            $240,000

b.            $266,000

c.             $260,000

d.            $250,000

2.        On the statement of net assets in liquidation at June 30, 2020, total liabilities are:

a.            $256,000

b.            $250,000

c.             $264,000

d.            $275,000
Answer #1

(1) On the statement of net assets in liquidation at June 30, 2020, total assets are: (c) 260,000

valuation of asset is as follows:-

  • Inventories- $55,000
  • Plant and equipment-$185,000
  • cash- (10,000+48,000+100,000+125,000+72,000) - (325,000+10,000) = 20,000

Total asset= 55,000 + 1,85,000+ 20,000 = 260,000

(2)  On the statement of net assets in liquidation at June 30, 2020, total liabilities are: (a) 256,000

creditors= 250,000

Liquidation cost= 6,000